What is the resistance a 1.60m long copper wire that is 0.100mm in diameter? A 90.0 long piece of carbon with a 1.00mm x 1.00mm square cross section?
Answer #1

Determine the cross-sectional area of the copper wire.

$$ \begin{aligned} A &=\pi\left(\frac{d}{2}\right)^{2} \\ &=\pi\left(\frac{0.100 \mathrm{~mm}\left(\frac{10^{-3} \mathrm{~m}}{1 \mathrm{~mm}}\right)}{2}\right)^{2} \\ &=7.854 \times 10^{-9} \mathrm{~m}^{2} \end{aligned} $$

The resistance of the copper wire is,

$$ \begin{aligned} R &=\frac{\rho L}{A} \\ &=\frac{\left(1.68 \times 10^{-8} \Omega \cdot \mathrm{m}\right)(1.60 \mathrm{~m})}{7.854 \times 10^{-9} \mathrm{~m}^{2}} \\ &=3.42 \Omega \end{aligned} $$

Here, \(\rho\) is the resistivity of the wire and \(L\) is the length of the wire.

(2)

Determine the cross-sectional area of the carbon wire.

$$ \begin{aligned} A &=\left(1.00 \mathrm{~mm}\left(\frac{10^{-3} \mathrm{~m}}{1 \mathrm{~mm}}\right)\right)^{2} \\ &=1.00 \times 10^{-6} \mathrm{~m}^{2} \end{aligned} $$

The resistance of the carbon wire is,

$$ \begin{aligned} R &=\frac{\rho L}{A} \\ &=\frac{\left(3.5 \times 10^{-5} \Omega \cdot \mathrm{m}\right)\left(90.0 \mathrm{~cm}\left(\frac{10^{-2} \mathrm{~m}}{1 \mathrm{~cm}}\right)\right)}{1.00 \times 10^{-6} \mathrm{~m}^{2}} \\ &=31.5 \Omega \end{aligned} $$